OpenStudy (anonymous):

Find the point on C where the curve 2x^2 - 3xy - y^3 = 1 has a horizontal tangent line? I found dy/dx = (4x-3y) / (3x+3y^2). Then I set 4x-3y = 0 and got y = 4/3x. After substituting y = 4/3x into the original equation, I got 2x^2 - 4x^2 - (4/3x)^3 = 1, which when simplified, is -2x^2 - (4/3x)^3 = 1. How do I solve for x in this case?

OpenStudy (nincompoop):

you're using implicit derivative?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

k

OpenStudy (nincompoop):

you plugged your y = 4/3x into all the y in the original equation properly?

OpenStudy (anonymous):

i think so. i checked three times. do you think i might have made a mistake in the process?

OpenStudy (nincompoop):

let me see

OpenStudy (nincompoop):

2x^2 - 3xy - y^3 = 1 2x^2 - 3x(4/3x) - (4/3x)^3 = 1 2x^2 - 12x/3x - (64/27x^3) = 1 -64/(27 x^3)+2 x^2-4 = 1

OpenStudy (anonymous):

i see where you made 3x (4/3x) = 12x/3x. but how do i find x after this?

OpenStudy (anonymous):

we have both x^3 and x^2 in the equation...

OpenStudy (nincompoop):

something's amiss

OpenStudy (nincompoop):

show me the implicit derivation just to make sure we're not solving for the wrong x

OpenStudy (anonymous):

you know what, i actually have an idea! i'm solving it right now. i'll show you the work when it's done. it might take a few minutes.

OpenStudy (nincompoop):

k

OpenStudy (anonymous):

ah... that didn't work out. i'll show you my work from finding the slope to substituting y = 4/3x into the original equation.

OpenStudy (nincompoop):

the x is not in the denominator

OpenStudy (nincompoop):

y = 4x/3

OpenStudy (nincompoop):

that changes the whole thing

OpenStudy (anonymous):

sorry my computer froze

OpenStudy (anonymous):

i'm following your work right now..

OpenStudy (nincompoop):

2x^2 - 3xy - y^3 = 1 2x^2 - 3x(4x/3) - (4x/3)^3 = 1 2x^2 - (12x^2)/3 - 64x^3/9 = 1 2x^2 - 4x^2 - (64x^3)/9 = 1 I think it is easy from here on

OpenStudy (anonymous):

make all denominators 9?

OpenStudy (nincompoop):

or remove the denominator by multiplying 9 on both sides of the equation

OpenStudy (anonymous):

yes i did that

OpenStudy (anonymous):

i got -18x^2 - 64x^3 = 9 and -2x^2 (9+32x) = 9

OpenStudy (anonymous):

and now i'm solving for x by doing -2x^2 = 9 and 9+32x=9

OpenStudy (anonymous):

x^2 = -9/2 ... i don't know what x will be when i put a radical sign over -9/2. for 9+32x = 9 i got x = 0...

OpenStudy (nincompoop):

why?

OpenStudy (anonymous):

what do you mean... why?

OpenStudy (nincompoop):

you factored it out ?

OpenStudy (anonymous):

ah.... you know what! i give up.... :'(

OpenStudy (anonymous):

i dont think i can solve this problem :'((

OpenStudy (nincompoop):

the factoring is the most tedious part of the problem

OpenStudy (anonymous):

what about you?

OpenStudy (nincompoop):

x = -3/32-3/32 (-1)^(2/3) (3/(259-16 sqrt(262)))^(1/3)+1/32 3^(2/3) (16sqrt(262)-259)^(1/3) or x = -3/32+3/32 (-3/(259-16 sqrt(262)))^(1/3)-1/32 (-1)^(1/3) 3^(2/3) (16 sqrt(262)-259)^(1/3) or x = -3/32-3/32 (3/(259-16 sqrt(262)))^(1/3)+1/32 (-3)^(2/3) (16 sqrt(262)-259)^(1/3)

OpenStudy (anonymous):

oh my gosh... what is that... haha

OpenStudy (nincompoop):

I don't even want to know

OpenStudy (anonymous):

lol..... :)

OpenStudy (anonymous):

thanks for your help! sooooo much!! i probably took too much of your precious time..

OpenStudy (anonymous):

i thnk it's time for us to move on with this problem..

OpenStudy (nincompoop):

I'll handle this later

OpenStudy (anonymous):

haha, you don't have to :) you already did a lot for me :)

OpenStudy (anonymous):

thanks again! i'm closing the question.

OpenStudy (nincompoop):

k